OpenStudy (anonymous):

fnd all solutions on the interval [0,2pi] for the equation 2- sinx= 2cos^2x

OpenStudy (cwrw238):

2 - sin x = 2 - 2 sin^2x 2sin^2x - sin x = 0 sin x ( 2sinx - 1) = 0 can you continue from here?

OpenStudy (anonymous):

i will still be lost from there. what do you do next?

OpenStudy (cwrw238):

since the product = 0 either sin x = 0 or 2sinx - 1 = 0 so sin x = 0 or sinx = 1/2 sin x = 0 gives x = 0 , pi, 2pi sin x = 1/2 gives x = pi/6 , 5pi/6

OpenStudy (anonymous):

okayy, thank you so much.

OpenStudy (cwrw238):

no probs